# Polynomial time auditing of statistical subgroup fairness for Gaussian data

### 27 Jan 2024 v1

@@ Abstract
We study the problem of auditing classifiers with the notion of statistical subgroup fairness. [1] has shown that the problem of auditing combinatorial subgroups fairness is as hard as agnostic learning. Essentially all work on remedying statistical measures of discrimination against subgroups assumes access to an oracle for this problem, despite the fact that no efficient algorithms are known for it. If we assume the data distribution is Gaussian, or even merely log-concave, then a recent line of work has discovered efficient agnostic learning algorithms for halfspaces. Unfortunately, the boosting-style reductions given by Kearns et al. required the agnostic learning algorithm to succeed on reweighted distributions that may not be log-concave, even if the original data distribution was. In this work, we give positive and negative results on auditing for the Gaussian distribution: On the positive side, we an alternative approach to leverage these advances in agnostic learning and thereby obtain the first polynomial-time approximation scheme (PTAS) for auditing nontrivial combinatorial subgroup fairness: we show how to audit statistical notions of fairness over homogeneous halfspace subgroups when the features are Gaussian. On the negative side, we find that under cryptographic assumptions, no polynomial-time algorithm can guarantee any nontrivial auditing, even under Gaussian feature distributions, for general halfspace subgroups.
@@ 1 INTRODUCTION
The deployment of decision rules obtained using machine learning has raised the risk that the rules may exhibit biases against historically marginalized communities. In particular, [1] raised the concern that these decision rules may be biased against sub-groups characterized by a combination of “protected” attributes. Since there is an exponential number of such subgroups, even detecting such statistical patterns of discrimination is a nontrivial computational problem; indeed, [1] showed that the problem of finding disadvantaged subgroups is equivalent to the problem of agnostic learning, which is believed to be intractable in general for albeit the simplest classes of sets. Essentially all work [1, 2, 3] on remedying statistical measures of discrimination against subgroups assumes access to an oracle for this problem, despite the fact that no efficient algorithms are known for it. In this work we are proposing a solution for the fairness auditing problem with provable guarantees of efficiency and correctness.
@@ 1.1 Background and Motivation
In the court case “DeGraffenreid v General Motors" [6], five Black women brought suit against General Motors for its discrimination against the group of Black women. Although no sex discrimination was revealed, the evidence showed that Black women hired after 1970 were discriminated against by the company’s seniority system. Such discrimination can be better demonstrated by an example shown in table 1. In particular, the hiring rate of a company could seemingly be fair in terms of gender of race alone but clearly discriminates against the subgroups of white men and black women. As a result, the court rejected the plaintiffs’ attempt to bring a suit not on behalf of Blacks or women, but specifically on behalf of Black women. In the ruling, the judge was specifically concerned about the proliferation of protected classes.
@@
Thereafter, [1] proposed more general notions of statistical fairness that require auditing over subgroups defined on simple combinations of data features. Specifically, such combinations of features can be any simple representations, such as conjunctions and halfspaces, which, however, can generate exponentially many subgroups. They also showed that the problem of auditing subgroups defined by such simple representation is as hard as “weak agnostic learning” [7, 8]. While the problem of weak agnostic learning is widely believed to be computationally intractable [8, 9], its hardness does not necessarily hold for specific distributions. Thus, it is natural for someone to think of attacking the auditing problem through distribution-specific agnostic learning approaches as agnostic learning is a much more extensively studied problem. However, it turns out there are still obstacles remaining for doing so.
@@ 1.2 Challenges of Auditing through Agnostic Learning
The main challenge that prevents us from applying existing agnostic learning techniques to perform auditing based on the reduction by [1] is that it is formulated in terms of weak agnostic learning, that is, finding classifiers with error rates that are nonnegligibly better than guessing, and correspondingly weak auditing guarantees. What we would like instead is for the unfairness certificate to be near-optimal, so that the unfairness it exhibits certifies a bound on the unfairness of the classifier for all subgroups. When we have guarantees for arbitrary distributions, “boosting” [10] enables us to obtain high accuracy from such weak learners. Unfortunately, these techniques require re-weighting the data examples after which the distribution-specific properties may no longer hold.
One might hope to dodge this issue by casting the problem of finding a harmed subgroup as a Mixed-Integer Program and using solvers that, though they lack polynomial-time guarantees, obtain adequate performance in practice. In such an approach, the failure of the solver to find a feasible solution to the optimization problem is taken as the proof that the classifier is fair. Unfortunately, these solvers owe their speed in part to a lack of soundness, both due to numerical issues [11] and the complexity of the heuristics used to prune the search [12, 13], and it remains an open question whether or not it is possible to obtain acceptable performance (using the various advanced techniques employed by commercial solvers) while retaining the guarantee that the solver correctly reports infeasibility [14]. In any case, the works by [1, 15] and [16] that empirically studied these approaches to obtaining fair classifiers used linear regression as a proxy for the agnostic learning or cost-sensitive classification subroutines. Unfortunately, these heuristics do not even provide in-principle guarantees.
In this paper, we will present an alternative auditing approach for subgroups determined by homogeneous halfspaces with provable guarantees.
@@ 1.3 Our Contribution
Our main contribution is a more careful analysis of the relationship between auditing and agnostic learning: Given a fixed positive classification rate, the harm (w.r.t. statistical parity) suffered by a subgroup is affinely related to the error rate of the subgroup indicator. Thus, a solution to the agnostic learning problem directly gives a harmed subgroup. It is worth noting that under a standard normal distribution, the subclass of halfspaces with a fixed positive classification rate is given by the halfspaces with unit normal vectors and the same threshold.
@@
Furthermore, following [17], we also obtain a lower bound on the unfairness detectable when auditing for halfspace subgroups under Gaussian distributions by reducing the problem of continuous Learning With Errors (cLWE) to auditing.
For our algorithmic results, we will present a general auditing framework given an oracle for (distribution-specific) agnostic learning. Also, we give a randomized PTAS auditing algorithm for subgroups determined by homogeneous halfspaces under Gaussian data by applying the method from [18].
@@
At first blush, the reliance on a (prima facie unverifiable) distributional assumption for the analysis of our auditing algorithm may seem to be at odds with our desire to certify the fairness of a classifier. Nevertheless, a line of recent works by [19] and [20] have shown that the properties of the data that are crucial to these algorithms for distribution-specific learning of halfspaces can be verified. Thus, these methods give a way of certifying fairness for families of nice distributions: so long as the data passes these tests and the audit reveals no subgroup that is significantly harmed, we may guarantee that the classifier is fair.
This paper will be organized as follow. Some necessary notations for our arguments are given in section 2. Then, we will present the main reduction and the hardness result in section 3. Section 4 will present out auditing framework as well as the distribution-specific PTAS algorithm. At last, we will discuss the limitation of our approach and our future work.
@@ 1.4 Related Work
[1] gave an game theory based approach for learning representations with subgroup fairness by assuming there exists an efficient oracle for auditing. A follow-up study [15] was later conducted to estimate their algorithm on real-world datasets. [3] showed a method of obtaining “multi-accurate” representations by assuming the existence of an efficient auditing oracle. Besides, [2] proposed a variant of statistical fairness called “multi-fairness,” which allows them to efficiently learn a multi-fair classifier with querying “relative fairness” of data pairs. As we discussed previously, the auditing oracles in these works were provided by using linear regression as a heuristic for the optimal halfspace, which does not provide guarantees. They also did not consider auditing for specific families of distributions. On the other hand, the works on agnostic learning for specific families of distributions, e.g., [21, 22, 18, 23, 24] do not consider how their techniques may be applied to the subgroup fairness auditing problem.
@@ 2 PRELIMINARIES
Further, the probability of an event under a distribution $\mathcal{D}$ is denoted by $\Pr_{\mathbf{x}\sim\mathcal{D}}\{\cdot\}$. Specifically, $\mathcal{N}(0,\mathrm{I})$ denotes a standard normal distribution, where $\mathrm{I}$ represents the identity matrix. For simplicity of notation, we may use $\mathcal{N},\mathcal{N}_{\sigma}$ instead of $\mathcal{N}(0,\mathrm{I}),\mathcal{N}(0,\sigma^{2}\mathrm{I})$ or even drop $\mathcal{D}$ and $\mathcal{N}$ from the subscript when it is clear from the context.
@@
Equation (1) is a straightforward way to quantify how much the positive classification rate within a subgroup deviates from that of the overall population. The goal of fairness auditing is to develop an “auditing algorithm" to efficiently find such a certificate $g\in\mathcal{G}$ for any $c\in\mathcal{C}$ with sample access to $\mathcal{D}$.
Definition 2.3 ([1] Auditing Algorithm).
Under the notion of Statistical Parity Subgroup Fairness, fix a collection of group indicators $\mathcal{G}$ over the protected features, and any $\delta,\gamma,\gamma^{\prime}\in(0,1)$ such that $\gamma^{\prime}\leq\gamma$. A $(\gamma,\gamma^{\prime})$-auditing algorithm for $\mathcal{G}$ with respect to distribution $\mathcal{D}$ is an algorithm $\mathcal{A}$ such that for any classifier $h$, when given access the joint distribution $(\mathcal{D},h(\mathcal{D}))$, $\mathcal{A}$ runs in time $\mathrm{poly}(1/\gamma^{\prime},\log(1/\delta))$, and with probability $1-\delta$, outputs a $\gamma^{\prime}$-unfair certificate for $h$ whenever $h$ is $\gamma$-unfair with respect to $\mathcal{D}$ and $\mathcal{G}$. If $h$ is $\gamma^{\prime}$-fair, $\mathcal{A}$ will output “fair".
Since our reduction involves the subclass of halfspaces with a “probabilistic lock”, we give the formal definition of it as follows.
Definition 2.4 (Halfspaces With Probabilistic Lock).
We use $\mathcal{H}^{d}$ to represent the collection of all the halfspaces in $\mathbb{R}^{d}$. Then, for any arbitrary distribution $\mathcal{D}$ over $\mathbb{R}^{d}$, we define the collection of all halfspaces with the same probability threshold $\mu$ as
@@
To state the hardness results, we denote $\mathbb{S}^{d-1}:=\{\mathbf{x}\in\mathbb{R}^{d}\ |\ \lVert{\mathbf{x}}\rVert_{2}=1\}$, $\mathbb{Z}_{q}:=\{0,1,\ldots,q-1\}$, $\mathbb{R}_{q}:=[0,q)$, and $\mod_{q}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{q}$ for $q\in\mathbb{N}$. Then we formally define the problem of “learning with errors" (LWE) [25], following [17]:
@@
  • Null hypothesis: While $\mathbf{x}\in\mathcal{D}_{sp}$, $\mathbf{x}$ and $y$ are independent, where $y$ is sampled uniformly at random on the support of its marginal distribution in alternative hypothesis.
@@
3 FAIRNESS AUDITING AND AGNOSTIC LEARNING
In this section, we describe our reduction from auditing to agnostic learning, which is also our main contribution. In addition, we give a lower bound for fairness auditing under Gaussian distributions.
@@ 3.1 Reduction from Auditing to Halfspace Learning
While learning from a representation class like $\mathcal{H}_{\mu}^{\mathcal{D}}$ may seems to be hard at a first glance, there are actually examples [23] of learning $\mathcal{H}_{\mu}^{\mathcal{D}}$ in agnostic setting under Gaussian data. Actually, under Gaussian distribution, it is even simpler to consider iteratively learning $\mathcal{H}_{\mu}^{\mathcal{D}}$ with fixed $\mu$, because fixing the threshold allows only optimizing the direction of normal vector due to fact 2.1 (rotational invariance), which resembles to the task of learning homogeneous halfspaces.
@@
$s.t.\quad$ $\Pr_{\mathbf{x}\in\mathcal{D}}\{h(\mathbf{x})=1\}=a+k(b-a)/n.$
@@
$\max_{k}\Pr\{h_{k}^{*}\}\left|d_{\mathcal{D}}(c,h_{k}^{*})\right|\geq\gamma^{*}-2(b-a)/n.$
@@
$\left|\Pr\{h^{*}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h^{*})\right|-\left|\Pr\{h^{\prime}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h^{\prime})\right|\leq$ $|\Pr\{h^{*}\}-\Pr\{h^{\prime}\}|+|\Pr\{h^{\prime}\cap c\}-\Pr\{h^{*}\cap c\}|$
$\overset{\mathrm{}}{\leq}$ $(b-a)/n+|\Pr\{h^{\prime}\cap c\cap\mathbf{dom}\}-\Pr\{h^{*}\cap c\cap\mathbf{dom}\}|$
$\leq$ $(b-a)/n+|\Pr\{\mathbf{x}\in\mathbf{dom}\}|$
$\leq$ $2(b-a)/n$ (6)
@@
$\Pr\{h\}\left|d_{\mathcal{D}}(c,h)\right|=$ $|\Pr\{h^{*}\}d_{\mathcal{D}}(c,h^{*})+(\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h(\mathbf{x})\})/2|$
$\geq$ $\gamma^{*}-\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h(\mathbf{x})\}\right|/2$
@@
$\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h_{\mathbf{v}}(\mathbf{x})\}\right|/2\leq\epsilon$
@@
$\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h_{\mathbf{u}}(\mathbf{x})\}\right|/2\leq\epsilon$
@@
3.2 The Hardness of Auditing Under Gaussian Data

### 29 Feb 2024 v2

# Distribution-Specific Auditing For Subgroup Fairness

@@ Abstract
We study the problem of auditing classifiers for statistical subgroup fairness. Kearns et al. [1] showed that the problem of auditing combinatorial subgroups fairness is as hard as agnostic learning. Essentially all work on remedying statistical measures of discrimination against subgroups assumes access to an oracle for this problem, despite the fact that no efficient algorithms are known for it. If we assume the data distribution is Gaussian, or even merely log-concave, then a recent line of work has discovered efficient agnostic learning algorithms for halfspaces. Unfortunately, the reduction of Kearns et al. was formulated in terms of weak, “distribution-free” learning, and thus did not establish a connection for families such as log-concave distributions. In this work, we give positive and negative results on auditing for Gaussian distributions: On the positive side, we present an alternative approach to leverage these advances in agnostic learning and thereby obtain the first polynomial-time approximation scheme (PTAS) for auditing nontrivial combinatorial subgroup fairness: we show how to audit statistical notions of fairness over homogeneous halfspace subgroups when the features are Gaussian. On the negative side, we find that under cryptographic assumptions, no polynomial-time algorithm can guarantee any nontrivial auditing, even under Gaussian feature distributions, for general halfspace subgroups.
@@ 1 INTRODUCTION
The deployment of decision rules obtained using machine learning has raised the risk that the rules may exhibit biases against historically marginalized communities. In particular, Kearns et al. [1] raised the concern that these decision rules may be biased against sub-groups characterized by a combination of “protected” attributes. Since there are an exponential number of such subgroups, even detecting such statistical patterns of discrimination is a nontrivial computational problem; indeed, Kearns et al. [1] showed that the problem of finding disadvantaged subgroups is equivalent to the problem of agnostic learning, which is believed to be intractable in general for all but the simplest classes of sets. Essentially all work [1, 2, 3] on remedying statistical measures of discrimination against subgroups assumes access to an oracle for this problem, despite the fact that no efficient algorithms are known for it. In this work we are proposing a solution for a variant of the fairness auditing problem with provable guarantees of efficiency and correctness, as well as some strong limitations on the extent to which these solutions can be extended to richer families of subgroups.
@@ 1.1 Background and Motivation
In the court case “DeGraffenreid v General Motors" [6], five Black women brought suit against General Motors for its discrimination against the group of Black women. Although no sex discrimination was revealed, the evidence showed that Black women hired after 1970 were discriminated against by the company’s seniority system. Such discrimination can be better demonstrated by an example shown in table 1. In particular, the hiring rate of a company could seemingly be fair in terms of gender of race alone but clearly discriminates against the subgroups of white men and black women. As a result, the court rejected the plaintiffs’ attempt to bring a suit not on behalf of Blacks or women, but specifically on behalf of Black women. In the ruling, in favor of the defendant, the judge was specifically concerned about the proliferation of protected classes.
@@
Thereafter, [1] proposed more general notions of statistical fairness that require auditing over subgroups defined on simple combinations of data features. Specifically, such combinations of features can be any simple representations, such as conjunctions and halfspaces, which, however, can generate exponentially many subgroups. They also showed that the problem of auditing subgroups defined by such simple representation is as hard as “weak agnostic learning” in the standard “distribution-free” setting [7, 8]. While the problem of distribution-free weak agnostic learning is widely believed to be computationally intractable [8, 9], its hardness does not necessarily hold for specific distribution families. Thus, it is natural to consider auditing using distribution-specific agnostic learning approaches as agnostic learning is a much more extensively studied problem. However, it turns out there are still obstacles remaining for doing so.
@@ 1.2 Challenges of Auditing through Agnostic Learning
The main challenge that prevents us from applying existing agnostic learning techniques to perform auditing based on the reduction by [1] is that it is formulated in terms of weak agnostic learning, that is, finding classifiers with error rates that are nonnegligibly better than guessing, and correspondingly weak auditing guarantees. In particular, the approximation guarantees we obtain for distribution-specific agnostic learning yield vacuous guarantees for weak learning. When we have guarantees for arbitrary distributions, “boosting” [10] enables us to obtain high accuracy from such weak learners. Unfortunately, these techniques require re-weighting the data examples after which the distribution-specific properties may no longer hold.
One might hope to dodge this issue by casting the problem of finding a harmed subgroup as a Mixed-Integer Program and using solvers that, though they lack polynomial-time guarantees, obtain adequate performance in practice. In such an approach, the failure of the solver to find a feasible solution to the optimization problem is taken as the proof that the classifier is fair. Unfortunately, these solvers owe their speed in part to a lack of soundness, both due to numerical issues [11] and the complexity of the heuristics used to prune the search [12, 13], and it remains a current research challenge to obtain acceptable performance (using the various advanced techniques employed by commercial solvers) while retaining the guarantee that the solver correctly reports infeasibility [14]. In any case, the works by [1, 15] and [16] that empirically studied these approaches to obtaining fair classifiers used linear regression as a proxy for the agnostic learning or cost-sensitive classification subroutines. Unfortunately, these heuristics do not even provide in-principle guarantees.
In this paper, we will show auditing general halfspace subgroups is hard even under Gaussian distribution and present an alternative auditing approach for subgroups determined by homogeneous halfspaces with provable guarantees.
@@ 1.3 Our Contribution
Our first contribution is a more careful analysis of the relationship between auditing and agnostic learning: Given a fixed positive classification rate, the harm (w.r.t. statistical parity) suffered by a subgroup is affinely related to the error rate of the subgroup indicator. Thus, a solution to the agnostic learning problem directly gives a harmed subgroup. It is worth noting that under a standard normal distribution, the subclass of halfspaces with a fixed positive classification rate is given by the halfspaces with unit normal vectors and the same threshold.
@@
Based on the reduction and a inspiration from Diakonikolas et al. [17], our second major contribution is a lower bound on the unfairness detectable when auditing for halfspace subgroups under Gaussian distributions by reducing the problem of continuous Learning With Errors (cLWE) to auditing. Our hardness results include both multiplicative and additive forms. More interestingly, we can further show that even “nonconstructive auditing” is hard, where we do not need to exhibit a discriminated subgroup for a failed audit.
For our algorithmic results, we will present a general auditing framework given an oracle for (distribution-specific) agnostic learning. Also, we give a randomized PTAS auditing algorithm for subgroups determined by homogeneous halfspaces under Gaussian data by applying the method from Diakonikolas et al. [18].
@@
At first blush, the reliance on a (prima facie unverifiable) distributional assumption for the analysis of our auditing algorithm may seem to be at odds with our desire to certify the fairness of a classifier. Nevertheless, a line of recent works by Rubinfeld and Vasilyan [19] and Gollakota et al. [20] have shown that the properties of the data that are crucial to these algorithms for distribution-specific learning of halfspaces can be verified. Thus, these methods give a way of certifying fairness for families of nice distributions: so long as the data passes these tests and the audit reveals no subgroup that is significantly harmed, we may guarantee that the classifier is fair.
This paper will be organized as follow. Some necessary background for our arguments are given in Section 2. We will present the main reduction from auditing to agnostic learning in Section 3. Then, we will show the hardness results in Section 4. Section 5 will present out auditing framework as well as the distribution-specific PTAS algorithm. At last, we will discuss the limitation of our approach and our future work.
@@ 1.4 Related Work
Many authors have considered the problem of ensuring fairness in classification, and Barocas et al. [21] give a good overview of the broader area. In particular, there are alternatives to the statistical, group-fairness notions we are considering, for example individual-level fairness as proposed by Dwork et al. [22], or based on causal modeling, such as the “counterfactual” fairness notion proposed by Kusner et al. [23]. We cannot do justice to the breadth of literature and philosophical issues here, and we strongly encourage the interested reader to consult Barocas et al. The group-fairness notions we consider have their roots in the game-theory-based approach of Kearns et al. [1] for learning representations with subgroup fairness by assuming there exists an efficient oracle for auditing. A follow-up study [15] evaluated their algorithm on real-world datasets. Hébert-Johnson et al. [3] showed a method of obtaining “multi-accurate” representations by assuming the existence of an efficient auditing oracle. Further, Kim et al. [2] proposed a variant of statistical fairness called “multi-fairness,” which allows them to efficiently learn a multi-fair classifier with querying “relative fairness” of data pairs. As we discussed previously, the auditing oracles in these works were provided by using linear regression as a heuristic for the optimal halfspace, which does not provide guarantees. They also did not consider auditing for specific families of distributions. On the other hand, the works on agnostic learning for specific families of distributions, e.g., [24, 25, 18, 26, 27] do not consider how their techniques may be applied to the subgroup fairness auditing problem.
@@ 2 PRELIMINARIES
Further, the probability of an event under a distribution $\mathcal{D}$ is denoted by $\Pr_{\mathbf{x}\sim\mathcal{D}}\{\cdot\}$. $\mathcal{N}(0,\mathrm{I})$ denotes a standard normal distribution, where $\mathrm{I}$ represents the identity matrix. For simplicity of notation, we may use $\mathcal{N},\mathcal{N}_{\sigma}$ instead of $\mathcal{N}(0,\mathrm{I}),\mathcal{N}(0,\sigma^{2}\mathrm{I})$ or even drop $\mathcal{D}$ and $\mathcal{N}$ from the subscript when it is clear from the context.
@@
Equation (1) is a straightforward way to quantify how much the positive classification rate within a subgroup deviates from that of the overall population. The weighting by the size of the group (i.e., $\Pr_{\mathbf{x}\sim\mathcal{D}}\{\mathbf{x}\in g\}$) is a concession to address the statistical issues that arise with estimating $d$ on small groups: we cannot escape that our empirical estimates are less accurate as the size shrinks. The goal of fairness auditing is to develop an “auditing algorithm" to efficiently find such a certificate $g\in\mathcal{G}$ for any $c\in\mathcal{C}$ with sample access to $\mathcal{D}$, formalized as follows.
Definition 2.3 (Constructive Auditing [1]).
Fix a collection of group indicators $\mathcal{G}$ over the protected features, and any $\delta,\gamma,\gamma^{\prime}\in(0,1)$ such that $\gamma^{\prime}\leq\gamma$. A constructive $(\gamma,\gamma^{\prime})$-auditing algorithm for $\mathcal{G}$ with respect to distribution $\mathcal{D}$ is an algorithm $\mathcal{A}$ such that for any classifier $h$, when given access the joint distribution $(\mathcal{D},h(\mathcal{D}))$, $\mathcal{A}$ runs in time $\mathrm{poly}(1/\gamma^{\prime},\log(1/\delta))$, and with probability $1-\delta$, outputs a $\gamma^{\prime}$-unfair certificate for $h$ whenever $h$ is $\gamma$-unfair with respect to $\mathcal{D}$ and $\mathcal{G}$. If $h$ is $\gamma^{\prime}$-fair, $\mathcal{A}$ will output “fair”.
Moreover, we will consider a more general type of auditing task, called “non-constructive auditing”, where the algorithms are only required to tell if a discriminated subgroup exists.
Definition 2.4 (Non-constructive Auditing).
Under the same setting as Definition 2.3, a non-constructive $(\gamma,\gamma^{\prime})$-auditing algorithm for $\mathcal{G}$ with respect to distribution $\mathcal{D}$ is an algorithm $\mathcal{A}$ such that for any classifier $h$, when given access the joint distribution $(\mathcal{D},h(\mathcal{D}))$, $\mathcal{A}$ runs in time $\mathrm{poly}(1/\gamma^{\prime},\log(1/\delta))$, and with probability $1-\delta$, claims $h$ is $\gamma^{\prime}$-unfair whenever $h$ is $\gamma$-unfair with respect to $\mathcal{D}$ and $\mathcal{G}$. If $h$ is $\gamma^{\prime}$-fair, $\mathcal{A}$ will output “fair”.
Since our reduction involves the subclass of halfspace subgroups of a fixed size, we give the formal definition of it as follows.
Definition 2.5 (Fixed-size Halfspaces).
We use $\mathcal{H}^{d}$ to represent the collection of all the halfspaces in $\mathbb{R}^{d}$. Then, for any arbitrary distribution $\mathcal{D}$ over $\mathbb{R}^{d}$, we define the collection of all halfspaces with the same (relative) density $\mu$ as
@@
To state the hardness results, we denote $\mathbb{S}^{d-1}:=\{\mathbf{x}\in\mathbb{R}^{d}\ |\ \lVert{\mathbf{x}}\rVert_{2}=1\}$, $\mathbb{Z}_{q}:=\{0,1,\ldots,q-1\}$, $\mathbb{R}_{q}:=[0,q)$, and $\mod_{q}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{q}$ by the unique translation by $q\mathbb{Z}^{d}$ for $q\in\mathbb{N}$. Then we formally define the problem of “learning with errors” (LWE) [28], following [17]:
@@
  • Null hypothesis: $\mathbf{x}\in\mathcal{D}_{sp}$, $y$ is sampled uniformly at random on the support of its marginal distribution in alternative hypothesis, independent of $\mathbf{x}$.
@@
3 FROM AUDITING TO AGNOSTIC LEARNING
In this section, we describe our reduction from auditing to agnostic learning. In addition, we give a lower bound for fairness auditing under Gaussian distributions.
@@ 3.1 Reduction from Auditing to Halfspace Learning
While learning from a representation class like $\mathcal{H}_{\mu}^{\mathcal{D}}$ may seems to be hard at a first glance, there are actually examples [26] of learning $\mathcal{H}_{\mu}^{\mathcal{D}}$ in an agnostic setting under Gaussian data.
@@
$s.t.\quad$ $\Pr_{\mathbf{x}\in\mathcal{D}}\{h(\mathbf{x})=1\}=a+\frac{k(b-a)}{n}.$
@@
$\max_{k}\Pr\{h_{k}^{*}\}\left|d_{\mathcal{D}}(c,h_{k}^{*})\right|\geq\gamma^{*}-\frac{2(b-a)}{n}.$
@@
$\left|\Pr\{h^{*}\}d_{\mathcal{D}}(c,h^{*})\right|-\left|\Pr\{h^{\prime}\}d_{\mathcal{D}}(c,h^{\prime})\right|\leq$ $|\Pr\{h^{*}\}-\Pr\{h^{\prime}\}|+|\Pr\{h^{\prime}\cap c\}-\Pr\{h^{*}\cap c\}|$
$\overset{\mathrm{}}{\leq}$ $\frac{(b-a)}{n}+|\Pr\{h^{\prime}\cap c\cap\mathbf{dom}\}-\Pr\{h^{*}\cap c\cap\mathbf{dom}\}|$
$\leq$ $\frac{(b-a)}{n}+|\Pr\{\mathbf{x}\in\mathbf{dom}\}|$
$\leq$ $\frac{2(b-a)}{n}$ (6)
@@
$\Pr\{h\}\left|d_{\mathcal{D}}(c,h)\right|=$ $|\Pr\{h^{*}\}d_{\mathcal{D}}(c,h^{*})+\frac{\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h(\mathbf{x})\}}{2}|$
$\geq$ $\gamma^{*}-\frac{\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h(\mathbf{x})\}\right|}{2}$
@@
$\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h_{\mathbf{v}}(\mathbf{x})\}\right|\leq 2\epsilon$
@@
$\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h_{\mathbf{u}}(\mathbf{x})\}\right|\leq 2\epsilon$
@@
4 The Hardness of Auditing Under Gaussian Data